OpenStudy (anonymous):

I WILL GIVE YOU A MEDEL IF YOU CAN ANSWER THIS QUESTION Becca's family has 4 gallons of milk. They drink 5/8 gallon per day. How long will their milk last? A. 2 1/2days B. 4 days C. 4 5/8days D. 6 2/5days

OpenStudy (anonymous):

♫let it go, let it go, ...♫

OpenStudy (anonymous):

@wolfe8 @mathmale :D

OpenStudy (anonymous):

D is the answer

OpenStudy (anonymous):

D

OpenStudy (wolfe8):

You will divide 4 with 5/8

OpenStudy (anonymous):



OpenStudy (mathmale):

Sometimes using units of measurement makes a problem (and its answer) a lot clearer. We know that the family consumes 5/8 gallon per day of milk. This can be written as \[\frac{ 5 gal }{ 1 day }\] This is called a "rate" (here, it's the milk consumption rate). The family starts out with 4 gallons of milk. We are to determine how long that milk will last. Obviously, the answer must be in days, not in gallons or consumption rates. Writing

OpenStudy (anonymous):

@mathmale Typo: \[\frac{\frac{5}{8}gal}{1 day}\] is the rate.

OpenStudy (anonymous):

\[4 gal \times \frac{1day}{\frac{5}{8}gal} = \frac{4gal \times 1 day}{\frac{5}{8}gal} = \frac{4}{\frac{5}{8}}days\]

OpenStudy (anonymous):

gallon units cancel when multiplying across

OpenStudy (mathmale):

Nice work, @osx! Glad you caught my typo.